# Agnostic Process Tomography

### v1

@@ 4.5 Low-degree unitaries
We use $\mathcal{L}_{n,d}$ to denote the class of $n$-qubit unitaries of degree $d$. Our first algorithm given in Algorithm 3 is inspired by the algorithm of [22]. It involves estimating every low-degree Pauli coefficient, resulting in a query complexity of $\tilde{\mathcal{O}}(n^{2d})$. Our second algorithm given in Algorithm 4 is the low-degree learning algorithm from [30], which uses the non-commutative Bohnenblust-Hille (BH) inequality of [116] to obtain a sample complexity independent of $n$. However, in the agnostic setting, Algorithm 4 does not achieve a sample complexity independent of $n$. From the guarantees we show in Theorem 4, Algorithm 4 is only efficient when the optimal error is at most inverse-exponential in the system size. In particular, when the unknown unitary is a member of the concept class, we recover the complexity of the realizable setting. In most other regimes for $\mathsf{opt}$, it is better to use Algorithm 3. We discuss Algorithm 4 to illustrate that straightforwardly modifying algorithms from the realizable setting does not always give good agnostic learning algorithms.
@@
1:Using $N$ queries to $U$, construct estimates $\alpha_{x}$ of the Pauli coefficients for all $x\in\mathcal{P}_{n}$ s.t. $|x|\leq d$.
@@
$N=\mathcal{O}\left(\frac{d\cdot 16^{d}n^{2d}\log(n/\delta)}{\epsilon^{2}}\right)$ (4.120)
@@
3:Construct $\chi_{c}=\{x\in\mathcal{P}_{n}:|\hat{U}_{x}^{\prime}|\geq c,|x|\leq d\}$.
@@
We start by estimating all $\mathcal{O}(n^{d}4^{d})$ Pauli coefficients of degree $d$. We use Lemma 11 $\mathcal{O}(n^{d}4^{d})$ times, to obtain estimates $\alpha_{x}$ such that with probability at least $1-\delta$,
$|\alpha_{x}-\hat{U}_{x}|\leq\epsilon_{1},\quad\forall x\in\mathcal{P}_{n}\text{ s.t. }|x|\leq d.$ (4.133)
@@
$N=\mathcal{O}\left(\frac{d\cdot 4^{d}n^{d}\log(n/\delta)}{\epsilon_{1}^{2}}\right).$ (4.134)
@@
$\leq\frac{n^{d}4^{d}\epsilon_{1}^{2}}{2}.$ (4.147)
The inequality follows because there are at most $n^{d}4^{d}$ $x$’s such that $|x|\leq d$. Thus, with probability at least $1-\delta$, we have
$d_{F}(U,V)\leq\mathsf{opt}(U,\mathcal{L}_{n,d})+\frac{\sqrt{n^{d}4^{d}}\epsilon_{1}}{\sqrt{2}}$ (4.148)
To obtain the 1-agnostic learning guarantee, we choose $\epsilon_{1}=\epsilon\sqrt{2}/\sqrt{n^{d}4^{d}}$. This gives us a query complexity
$N=\mathcal{O}\left(\frac{d\cdot 16^{d}n^{2d}\log(n/\delta)}{\epsilon^{2}}\right)$ (4.149)
@@
Now, we will use the non-commutative BH inequality of [116] to prove Theorem 4.
Lemma 17 (Non-commutative BH inequality; Theorem 1.2 in [116]).
@@
Here, we consider the learning algorithm of [30], and our analysis is similar to theirs. The algorithm is sketched in Algorithm 4.
Let $c$ be a constant which we will specify later. In Step 1 of Algorithm 4, we prepare $N_{1}$ copies of the Choi state $|v(U)\rangle$ and measure each of them in the basis $\{|v(\sigma_{x})\rangle\}_{x}$, where again we specify $N_{1}$ later. This allows us to construct empirical estimates $\{|\hat{U}^{\prime}_{x}|^{2}\}_{x}$ by sampling from the distribution $\{|\hat{U}_{x}|^{2}\}_{x}$. We wish to construct estimates such that, with probability at least $21-\delta/$ 2,
$||\hat{U}^{\prime}_{x}|^{2}-|\hat{U}_{x}|^{2}|\leq c^{2},\quad\forall x\in\mathcal{P}_{n}.$ (4.152)
@@
$||\hat{U}^{\prime}_{x}|-|\hat{U}_{x}||\leq c,\quad\forall x\in\mathcal{P}_{n}.$ (4.154)
@@
$\chi_{c}\triangleq\{x\in\mathcal{P}_{n}:|\hat{U}^{\prime}_{x}|\geq c,|x|\leq d\}$ (4.155)
@@
where we use Equation 4.154 and the definition of $\chi_{c}$. In Step 3 of Algorithm 4, we want to construct estimates $\alpha_{x}$ for the Pauli coefficients of all strings $x\in\chi_{c}$, such that with probability at least $21-\delta/$ 2,
$|\alpha_{x}-\hat{U}_{x}|\leq c\epsilon,\quad\forall x\in\chi_{c}$ (4.158)
@@
It remains to bound $d_{F}(\tilde{U}_{d},V)$, for which we follow the arguments of [30]. Denote by $W_{x}$ the Pauli coefficients of $\tilde{U}_{d}$. Then, we have
@@
$\|\tilde{U}_{d}\|_{\infty}\leq\min\left[n^{d/2}2^{d},1+2^{(n+1)/2}\cdot\mathsf{opt}(U,\mathcal{L}_{n,d})\right].$ (4.168)
@@
$\leq\sqrt{n^{d}4^{d}}\sqrt{\sum_{|x|\leq d}|\hat{U}_{x}|^{2}}$ (4.177)
$=n^{d/2}2^{d}\sqrt{w^{\leq d}(U)}\leq n^{d/2}2^{d},$ (4.178)
@@
$w^{\leq d}(W)=1\quad\text{ and }\quad d_{F}(U,W)=\mathsf{opt}(U,\mathcal{L}_{n,d}).$ (4.179)
@@ 5.1 Bounded-gate circuits
As in Section 4.1, the time complexity in the proposition depends on the complexity of computing the inner products of pairs of states. This is again due to the use of Clifford classical shadows [111], which is not computationally efficient for arbitrary observables (see Lemma 13). We do not analyze the time complexity in detail because the scaling with $|\mathcal{C}|$ is already computationally inefficient for most classes of interest.
@@
The algorithm and analysis are similar to the proof of Proposition 1. Namely, we use Clifford classical shadows [111] to estimate the distance $\mathrm{d_{avg}}(\mathcal{U}_{i},\mathcal{E})$ between the unknown channel $\mathcal{E}$ and all unitaries $U_{i}\in\mathcal{C}$, where again $\mathcal{U}_{i}(\rho)=U_{i}\rho U_{i}^{\dagger}$. Then, our algorithm outputs the one with the minimal estimated distance.
@@
[…]
Note that because our unknown channel may not be a unitary, the quantity estimated is not the root mean squared trace distance but instead the average infidelity. This is clear from the proof in Section C.2(a) of [27]. For any measure $\mu$ over input states, we denote the average infidelity by
@@
Now, by Equation (C.8) of [27], using
@@
$|\overline{\mathrm{F}}_{P}(\mathcal{U}_{i},\mathcal{E})-(1-\hat{o}_{i})|\leq\epsilon^{\prime},\quad\forall i\in[|\mathcal{C}|]$ (5.7)
with probability at least $1-\delta$ for some $\epsilon^{\prime}>0$. Then, we can select $i^{\star}=\arg\min_{i}(1-\hat{o}_{i})$ and output $U_{i^{\star}}$, i.e., the $U_{i}\in\mathcal{C}$ with the smallest estimated distance from $\mathcal{E}$ with respect to $\overline{\mathrm{F}}_{P}$. To relate this back to $\mathrm{d_{avg}}$, we recall that $P$ forms a $3$-design [137, 138, 139, 140, 141], implying that the average infidelity over $P$ is the same as that over the Haar measure. This gives us the following relation between the average distances:
$d_{P}(\mathcal{U}_{i},\mathcal{E})^{2}$ $=\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim P}[\mathrm{d_{tr}}((\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi})))^{2}]$ (5.8)
$\leq\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim P}[1-F(\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]$ (5.9)
$=\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim\mu_{S}}[1-F(\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]$ (5.10)
$\leq\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim\mu_{S}}[\mathrm{d_{tr}}(\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]$ (5.11)
$\leq\sqrt{\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim\mu_{S}}[\mathrm{d_{tr}}(\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))^{2}]}$ (5.12)
@@
$\mathrm{d_{avg}}(\mathcal{U}_{i^{\star}},\mathcal{E})^{2}$ $=\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim\mu_{S}}[\mathrm{d_{tr}}(\mathcal{U}_{i^{\star}}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))^{2}]$ (5.14)
$\leq\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim\mu_{S}}[1-F(\mathcal{U}_{i^{\star}}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]$ (5.15)
$=\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim P}[1-F(\mathcal{U}_{i^{\star}}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]$ (5.16)
@@
$\leq\min_{i}\left(\mathop{\mathbf{E}}\limits_{|\psi\rangle\sim P}[\mathrm{d_{tr}}(\mathcal{U}_{i}(\outerproduct{\psi}{\psi}),\mathcal{E}(\outerproduct{\psi}{\psi}))]\right)+2\epsilon^{\prime}$ (5.20)
@@
$V^{*}\triangleq\operatorname*{arg\,min}_{V\in\mathcal{C}}\mathrm{d_{avg}}(\mathcal{E},\mathcal{V})\quad\text{ and }\quad\mathsf{opt}(\mathcal{E},\mathcal{C})=\mathrm{d_{avg}}(\mathcal{E},\mathcal{V}^{*}).$ (5.31)
@@ 5.2 APT from agnostic state tomography
Let $01>\epsilon,\delta>$ 0. Suppose there exists a learning algorithm for (proper) $(1,\epsilon^{2},\delta)$-agnostic state tomography with respect to fidelity $F$ for some class $\mathcal{D}$ of pure states. Let $\mathcal{C}$ be a concept class of unitaries. Suppose that |$\ket{v(U)}\in\mathcal{D}$ for all $U\in\mathcal{C}$. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
As in Section 4.2, our algorithm simply applies agnostic state tomography to the Choi state of the unknown process, inspired by [87]. We also note that we can trivially obtain $(1,\epsilon,\delta)$-agnostic process tomography with respect to the entanglement infidelity from $(1,\epsilon,\delta)$-agnostic state tomography (notably, without the quadratic blowup with respect to $\epsilon$) under the same compatibility condition. Recall that we defined the entanglement infidelity in Equation 4.30, and this metric has been widely considered in (realizable) unitary estimation [87, 6, 88, 89, 90, 91, 92, 93].
@@
The proof is similar to that of Proposition 2. Let $\Phi$ denote the unknown unitary. The algorithm is simple: first, prepare the Choi state $v(\Phi)$ of $\Phi$ and then apply the $(1,\epsilon^{2},\delta)$-agnostic state tomography algorithm to $v(\Phi)$. Note that the known agnostic state tomography algorithms work for mixed state inputs [42, 43]. The output of our algorithm is the density matrix of the pure state output by the agnostic state tomography algorithm.
The agnostic state tomography algorithm outputs some state |$\ket{\psi}$ such that
$F(v(\Phi),\outerproduct{\psi}{\psi})\geq\max_{\ket{\phi}\in\mathcal{D}}F(v(\Phi),\outerproduct{\phi}{\phi})-\epsilon^{2},$ (5.40)
@@
$\mathrm{d_{tr}}(v(\Phi),\outerproduct{\psi}{\psi})^{2}\leq\min_{\ket{\phi}\in\mathcal{D}}\mathrm{d_{tr}}(v(\Phi),\outerproduct{\phi}{\phi})+\epsilon^{2}.$ (5.42)
We can relate this to the Frobenius norm using a result from [145]. Namely, they prove that
$\frac{1}{2}\norm{\rho-\sigma}_{F}^{2}\leq\mathrm{d_{tr}}(\rho,\sigma)^{2}\leq\frac{\mathrm{rank}(\rho)\mathrm{rank}(\sigma)}{\mathrm{rank}(\rho)+\mathrm{rank}(\sigma)}\norm{\rho-\sigma}_{F}^{2}.$ (5.43)
@@
Since |$\outerproduct{\phi}{\phi}$ | is pure, then $\mathrm{rank}(\outerproduct{\phi}{\phi})=1$ so that the term in the square root can be bounded by $1$. Moreover, by definition of the normalized Frobenius distance for operators (Definition 9),
$d_{F}(v(\Phi),\outerproduct{\psi}{\psi})=\frac{1}{2^{n}\sqrt{2}}\norm{v(\Phi)-\outerproduct{\psi}{\psi}}_{F}$ (5.45)
since $v(\Phi),\outerproduct{\psi}{\psi}$ | are on $2n$ qubits. Thus, rewriting the above inequality in terms of the normalized Frobenius distance, we have
$4^{n}d_{F}(v(\Phi),\outerproduct{\psi}{\psi})^{2}\leq\sqrt{2}\cdot 2^{n}\min_{\ket{\phi}\in\mathcal{D}}d_{F}(v(\Phi),\outerproduct{\phi}{\phi})+\epsilon^{2}.$ (5.46)
@@
$d_{F}(\Phi,\Psi)$ $=\frac{1}{2^{n}\sqrt{2}}\norm{\Phi-\Psi}_{F}$ (5.48)
$=\frac{1}{2^{n}\sqrt{2}}\norm{J(\Phi)-J(\Psi)}_{F}$ (5.49)
$=\frac{1}{2}\norm{v(\Phi)-v(\Psi)}_{F}$ (5.50)
@@
This gives a fairly general result for converting an agnostic state tomography algorithm with respect to $\mathcal{D}$ into an improper agnostic process tomography algorithm with respect to $\mathcal{C}$ when |$\ket{v(U)}\in\mathcal{D}$ for all $U\in\mathcal{C}$. However, our result is weaker than that of Proposition 2, as we have a dependence on $\sqrt{\mathsf{opt}}$ in the error bound, and the algorithm is no longer $1$-agnostic but $2^{1/4}$-agnostic.
As remarked earlier, the compatibility condition holds for many classes of unitaries for which agnostic state tomography algorithms already exist with respect to a suitable class. In particular, there exist agnostic state tomography algorithms for stabilizer states, states with high stabilizer dimension, and discrete product states [42, 43]. Note that the corresponding classes of unitaries, i.e., Clifford circuits [143], Clifford + T circuits, and unitaries consisting of a tensor product of single-qubit gates, respectively, all satisfy the required property of Proposition 4. Thus, similarly to Section 4.2, we can directly apply Proposition 2 with the state-of-the-art sample/time complexity bounds for agnostic state tomography to obtain improper APT algorithms for the respective unitary classes. These are each direct corollaries of Corollary 6.3, Theorem 7.1, Corollary 8.3, and in [43].
Corollary 9 (Improper APT of Clifford Circuits; via Corollary 6.3 in [43]).
Let $01>\epsilon,\delta>$ 0. Let $\mathcal{C}$ be the class of $n$-qubit Clifford unitaries. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
Corollary 10 (Improper APT of Clifford + T Circuits; via Theorem 7.1 in [43]).
Let $01>\epsilon,\delta>$ 0 and $t\geq 0$. Let $\mathcal{C}$ be the class of $n$-qubit unitaries consisting of Clifford gates and $t$ T gates. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
Corollary 11 (Improper APT of Product Circuits; via Corollary 8.3 in [43]).
Let $\mathcal{K}$ be a set of single-qubit gates for which $|\innerproduct{v(U_{1})}{v(U_{2})}|^{2}\leq 1-\mu$ for any distinct $U_{1},U_{2}\in\mathcal{K}$. Let $\mathcal{C}=\mathcal{K}^{\otimes n}$. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
Corollary 12 (Improper APT of Clifford Product Circuits; via Corollary 9.3 in [43]).
Let $01>\epsilon,\delta>$ 0. Let $\mathcal{C}$ be the class of tensor products of single-qubit Clifford gates. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that

### v2

@@ 4.5 Low-degree unitaries
We use $\mathcal{L}_{n,d}$ to denote the class of $n$-qubit unitaries of degree $d$. We note that there are ${n\choose d}3^{d}\leq n^{d}3^{d}$ Pauli strings of degree exactly $d$, and thus $\mathcal{O}(n^{d}3^{d})$ strings of degree at most $d$.
Our first algorithm given in Algorithm 3 is inspired by the algorithm of [nadimpalli2024pauli]. It involves estimating every low-degree Pauli coefficient, resulting in a query complexity of $\tilde{\mathcal{O}}(n^{2d})$. Our second algorithm given in Algorithm 4 is the low-degree learning algorithm from [arunachalam2024learning], which uses the non-commutative Bohnenblust-Hille (BH) inequality of [volberg2024noncommutative] to obtain a sample complexity independent of $n$. However, in the agnostic setting, Algorithm 4 does not achieve a sample complexity independent of $n$. From the guarantees we show in Theorem 4, Algorithm 4 is only efficient when the optimal error is at most inverse-exponential in the system size. In particular, when the unknown unitary is a member of the concept class, we recover the complexity of the realizable setting. In most other regimes for $\mathsf{opt}$, it is better to use Algorithm 3. We discuss Algorithm 4 to illustrate that straightforwardly modifying algorithms from the realizable setting does not always give good agnostic learning algorithms.
@@
1:Using NN queries to $U$, construct estimates $\alpha_{x}$ of the Pauli coefficients for all $x\in\mathcal{P}_{n}$ s.t. $|x|\leq d$.
@@
$N=\mathcal{O}\left(\frac{d\cdot 9^{d}n^{2d}\log(n/\delta)}{\epsilon^{2}}\right)$ (4.120)
@@
3:Construct $\chi_{c}=\{x\in\mathcal{P}_{n}\mathrel{\mathop{\ordinarycolon}}|\hat{U}_{x}^{\prime}|\geq c,|x|\leq d\}$.
@@
We start by estimating all $\mathcal{O}(n^{d}3^{d})$ Pauli coefficients of degree $d$. We use Lemma 11 $\mathcal{O}(n^{d}3^{d})$ times, to obtain estimates $\alpha_{x}$ such that with probability at least $1-\delta$,
$|\alpha_{x}-\hat{U}_{x}|\leq\epsilon_{1},\qquad\forall x\in\mathcal{P}_{n}\text{ s.t. }|x|\leq d.$ (4.133)
@@
$N=\mathcal{O}\left(\frac{d\cdot 3^{d}n^{d}\log(n/\delta)}{\epsilon_{1}^{2}}\right).$ (4.134)
@@
$\leq\mathcal{O}\left(n^{d}3^{d}\epsilon_{1}^{2}\right).$ (4.147)
The inequality follows because there are at most $\mathcal{O}(n^{d}3^{d})$ Pauli strings $x$ such that $|x|\leq d$. Thus, with probability at least $1-\delta$, we have
$d_{F}(U,V)\leq\mathsf{opt}(U,\mathcal{L}_{n,d})+\mathcal{O}\left(\sqrt{n^{d}3^{d}}\epsilon_{1}\right).$ (4.148)
To obtain the 1-agnostic learning guarantee, we choose $\epsilon_{1}=\mathcal{O}(\epsilon/\sqrt{n^{d}3^{d}})$. This gives us a query complexity
$N=\mathcal{O}\left(\frac{d\cdot 9^{d}n^{2d}\log(n/\delta)}{\epsilon^{2}}\right)$ (4.149)
@@
Now, we will use the non-commutative BH inequality of [volberg2024noncommutative] to prove Theorem 4.
Lemma 17 (Non-commutative BH inequality; Theorem 1.2 in [volberg2024noncommutative]).
@@
Here, we consider the learning algorithm of [arunachalam2024learning], and our analysis is similar to theirs. The algorithm is sketched in Algorithm 4.
Let $c$ be a constant which we will specify later. In Step 1 of Algorithm 4, we prepare $N_{1}$ copies of the Choi state $|v(U)\rangle$ and measure each of them in the basis $\{|v(\sigma_{x})\rangle\}_{x}$, where again we specify $N_{1}$ later. This allows us to construct empirical estimates $\{|\hat{U}^{\prime}_{x}|^{2}\}_{x}$ by sampling from the distribution $\{|\hat{U}_{x}|^{2}\}_{x}$. We wish to construct estimates such that, with probability at least $1-\delta/2$,
$||\hat{U}^{\prime}_{x}|^{2}-|\hat{U}_{x}|^{2}|\leq c^{2},\qquad\forall x\in\mathcal{P}_{n}.$ (4.152)
@@
$||\hat{U}^{\prime}_{x}|-|\hat{U}_{x}||\leq c,\qquad\forall x\in\mathcal{P}_{n}.$ (4.154)
@@
$\chi_{c}\triangleq\{x\in\mathcal{P}_{n}\mathrel{\mathop{\ordinarycolon}}|\hat{U}^{\prime}_{x}|\geq c,|x|\leq d\}$ (4.155)
@@
where we use Equation 4.154 and the definition of $\chi_{c}$. In Step 3 of Algorithm 4, we want to construct estimates $\alpha_{x}$ for the Pauli coefficients of all strings $x\in\chi_{c}$, such that with probability at least $1-\delta/2$,
$|\alpha_{x}-\hat{U}_{x}|\leq c\epsilon,\qquad\forall x\in\chi_{c}$ (4.158)
@@
It remains to bound $d_{F}(\tilde{U}_{d},V)$, for which we follow the arguments of [arunachalam2024learning]. Denote by $W_{x}$ the Pauli coefficients of $\tilde{U}_{d}$. Then, we have
@@
$\|\tilde{U}_{d}\|_{\infty}\leq\min\left[n^{d/2}3^{d/2},1+2^{(n+1)/2}\cdot\mathsf{opt}(U,\mathcal{L}_{n,d})\right].$ (4.168)
@@
$\leq\sqrt{n^{d}3^{d}}\sqrt{\sum_{|x|\leq d}|\hat{U}_{x}|^{2}}$ (4.177)
$=n^{d/2}3^{d/2}\sqrt{w^{\leq d}(U)}$ (4.178)
$\leq n^{d/2}3^{d/2},$ (4.179)
@@
$w^{\leq d}(W)=1\qquad\text{ and }\qquad d_{F}(U,W)=\mathsf{opt}(U,\mathcal{L}_{n,d}).$ (4.180)
@@ 5.1 Bounded-gate circuits
As in Section 4.1, the time complexity in the proposition depends on the complexity of computing the inner products of pairs of states. This is again due to the use of Clifford classical shadows [huang2020predicting], which is not computationally efficient for arbitrary observables (see Lemma 13). We do not analyze the time complexity in detail because the scaling with $|\mathcal{C}|$ is already computationally inefficient for most classes of interest.
@@
The algorithm and analysis are similar to the proof of Proposition 1. Namely, we use Clifford classical shadows [huang2020predicting] to estimate the distance $\mathrm{d_{avg}}(\mathcal{U}_{i},\mathcal{E})$ between the unknown channel $\mathcal{E}$ and all unitaries $U_{i}\in\mathcal{C}$, where again $\mathcal{U}_{i}(\rho)=U_{i}\rho U_{i}^{\dagger}$. Then, our algorithm outputs the one with the minimal estimated distance.
@@
$P=\mathrm{Uniform}(\{|\psi\rangle=U|0\rangle,U\in\mathrm{Cl}(n)\}).$ (5.4)
Note that because our unknown channel may not be a unitary, the quantity estimated is not the root mean squared trace distance but instead the average infidelity. This is clear from the proof in Section C.2(a) of [zhao2023learning]. For any measure $\mu$ over input states, we denote the average infidelity by
@@
Now, by Equation (C.8) of [zhao2023learning], using
@@
$|\overline{\mathrm{F}}_{P}(\mathcal{U}_{i},\mathcal{E})-(1-\hat{o}_{i})|\leq\epsilon^{\prime},\qquad\forall i\in[|\mathcal{C}|]$ (5.7)
with probability at least $1-\delta$ for some $\epsilon^{\prime}>0$. Then, we can select $i^{\star}=\arg\min_{i}(1-\hat{o}_{i})$ and output $U_{i^{\star}}$, i.e., the $U_{i}\in\mathcal{C}$ with the smallest estimated distance from $\mathcal{E}$ with respect to $\overline{\mathrm{F}}_{P}$. To relate this back to $\mathrm{d_{avg}}$, we recall that $P$ forms a $3$-design [kueng2015qubit, webb2015clifford, zhu2017multiqubit], implying that the average infidelity over $P$ is the same as that over the Haar measure. This gives us the following relation between the average distances:
[…]
@@
[…]
@@
[…]
@@
$V^{*}\triangleq\operatorname*{arg\,min}_{V\in\mathcal{C}}\mathrm{d_{avg}}(\mathcal{E},\mathcal{V})\qquad\text{ and }\qquad\mathsf{opt}(\mathcal{E},\mathcal{C})=\mathrm{d_{avg}}(\mathcal{E},\mathcal{V}^{*}).$ (5.31)
@@ 5.2 APT from agnostic state tomography
[…]
@@
As in Section 4.2, our algorithm simply applies agnostic state tomography to the Choi state of the unknown process, inspired by [leung2000towards]. We also note that we can trivially obtain $(1,\epsilon,\delta)$-agnostic process tomography with respect to the entanglement infidelity from $(1,\epsilon,\delta)$-agnostic state tomography (notably, without the quadratic blowup with respect to $\epsilon$) under the same compatibility condition. Recall that we defined the entanglement infidelity in Equation 4.30, and this metric has been widely considered in (realizable) unitary estimation [leung2000towards, haah2023query, acin2001optimal, peres2002covariant, hayashi2006parallel, chiribella2005optimal, kahn2007fast, yang2020optimal].
@@
The proof is similar to that of Proposition 2. Let $\Phi$ denote the unknown unitary. The algorithm is simple: first, prepare the Choi state $v(\Phi)$ of $\Phi$ and then apply the $(1,\epsilon^{2},\delta)$-agnostic state tomography algorithm to $v(\Phi)$. Note that the known agnostic state tomography algorithms work for mixed state inputs [grewal2024agnostic, chen2024stabilizer]. The output of our algorithm is the density matrix of the pure state output by the agnostic state tomography algorithm.
[…]
@@
[…]
We can relate this to the Frobenius norm using a result from [coles2019strong]. Namely, they prove that
$\frac{1}{2}\mathinner{\!\left\lVert\rho-\sigma\right\rVert}_{F}^{2}\leq\mathrm{d_{tr}}(\rho,\sigma)^{2}\leq\frac{\mathrm{rank}(\rho)\mathrm{rank}(\sigma)}{\mathrm{rank}(\rho)+\mathrm{rank}(\sigma)}\mathinner{\!\left\lVert\rho-\sigma\right\rVert}_{F}^{2}.$ (5.43)
@@
[…]
since $|v(\Phi),\outerproduct{\psi}{\psi}$ are on $2n$ qubits. Thus, rewriting the above inequality in terms of the normalized Frobenius distance, we have
[…]
@@
$d_{F}(\Phi,\Psi)$ $=\frac{1}{2^{n}\sqrt{2}}\mathinner{\!\left\lVert\Phi-\Psi\right\rVert}_{F}$ (5.48)
$=\frac{1}{2^{n}\sqrt{2}}\mathinner{\!\left\lVert J(\Phi)-J(\Psi)\right\rVert}_{F}$ (5.49)
$=\frac{1}{2}\mathinner{\!\left\lVert v(\Phi)-v(\Psi)\right\rVert}_{F}$ (5.50)
@@
[…]
As remarked earlier, the compatibility condition holds for many classes of unitaries for which agnostic state tomography algorithms already exist with respect to a suitable class. In particular, there exist agnostic state tomography algorithms for stabilizer states, states with high stabilizer dimension, and discrete product states [grewal2024agnostic, chen2024stabilizer]. Note that the corresponding classes of unitaries, i.e., Clifford circuits [gottesman1997stabilizer], Clifford + T circuits, and unitaries consisting of a tensor product of single-qubit gates, respectively, all satisfy the required property of Proposition 4. Thus, similarly to Section 4.2, we can directly apply Proposition 2 with the state-of-the-art sample/time complexity bounds for agnostic state tomography to obtain improper APT algorithms for the respective unitary classes. These are each direct corollaries of Corollary 6.3, Theorem 7.1, Corollary 8.3, and in [chen2024stabilizer].
Corollary 9 (Improper APT of Clifford Circuits; via Corollary 6.3 in [chen2024stabilizer]).
Let $1>\epsilon,\delta>0$. Let $\mathcal{C}$ be the class of $n$-qubit Clifford unitaries. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
Corollary 10 (Improper APT of Clifford + T Circuits; via Theorem 7.1 in [chen2024stabilizer]).
Let $1>\epsilon,\delta>0$ and $t\geq 0$. Let $\mathcal{C}$ be the class of $n$-qubit unitaries consisting of Clifford gates and $t$ T gates. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that
@@
Corollary 11 (Improper APT of Product Circuits; via Corollary 8.3 in [chen2024stabilizer]).
[…]
@@
Corollary 12 (Improper APT of Clifford Product Circuits; via Corollary 9.3 in [chen2024stabilizer]).
Let $1>\epsilon,\delta>0$. Let $\mathcal{C}$ be the class of tensor products of single-qubit Clifford gates. Given query access to an unknown channel $\Phi$, there exists an improper learning algorithm outputting a hypothesis $\Phi^{\prime}$ such that